OpenStudy (anonymous):

Write a recursive formula for this sequence: 16, 40, 100, 250..... Use parenthesis for subscripts. For example, ^an can be written a(n)

OpenStudy (anonymous):

firstly try to find a pattern for \(a_0=16\) , \(a_1=40\) and \(a_2=100\)

OpenStudy (anonymous):

which states \(a_3\) in terms of \(a_1\) and \(a_0\)

OpenStudy (anonymous):

for example i see \[100=\frac{40}{2}+5*16\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do i find the formula

OpenStudy (anonymous):

well as u can see \[a_2=\frac{a_1}{2}+5a_0\]

OpenStudy (anonymous):

sorry forgot to mention some notations let \(a_0=16\) , \(a_1=40\) , \(a_2=100\) and \(a_3=250\)

OpenStudy (anonymous):

\[a_2=\frac{a_1}{2}+5a_0\]\[a_3=\frac{a_2}{2}+5a_1\]now u can generalize it

OpenStudy (anonymous):

do cross multiply

OpenStudy (anonymous):

no just generalize it for \(a_n\) in terms of \(a_{n-1}\) and \(a_{n-2}\) before that can u tell me what is \(a_4\) ?

OpenStudy (anonymous):

250

OpenStudy (anonymous):

thats \(a_3\) i ask u to find \(a_4\)

OpenStudy (anonymous):

100

OpenStudy (anonymous):

but how do i write these in a formula

OpenStudy (anonymous):

\[\huge a_n=\frac{a_{n-1}}{2}+5 a_{n-2}\] for \[\huge n \ge 2\]

OpenStudy (anonymous):

is that the formula